# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 6.2. An application of Rogawski’s results
    is the same as the dimension of the space $V^{2}(\pi_{f}))$ of Case 5 of [30, pp. 91-92], which according to loc. cit. is equal to 1, by our choice of the archimedean component of $\pi(\phi)$. We obtain then the desired Hecke submodule. Finally, to get the assertion on the type of $H^{2}(\phi)_{!}$, we apply the following two facts:
    (a) if $r$ is the weight of the pure Hodge structure on $V_{\lambda}$ (so that, by (22), $r=k$ in our case), then $H^{2}_{!}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ has weight $r+2$. Its Hodge substructure $H^{2}(\pi(\phi)_{f})$ has thus the same weight;
    (b) the Hodge types of $H^{2}(\pi(\phi)_{f})$ are those $(p,q)$’s appearing in the $(p,q)$-decomposition of the space $H^{2}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi(\phi)_{\infty}\otimes V_{k,{\mathbb{C}}})$, which is in turn determined by the $(p,q)$-decomposition of the discrete series representation $\pi(\phi)_{\infty}$. Now, by Remark 6.7 the discrete series representation $\pi(\phi)_{\infty}$ coincides with $\pi^{+}$, which is holomorphic by ([7, 1.3]).
@@
Suppose the hypotheses of Proposition 6.8 hold. For any small enough level $K$, the simple, level $K$ Hecke module $\pi^{K}(\phi)_{f}$ corresponds to a minimal ideal in the level $K$ Hecke algebra. Take an idempotent generator $e_{\pi(\phi)}$ of this ideal. By a standard argument (cfr. for example [35, Prop. 5.4]) we see that the Hodge structure $H^{2}(\phi)_{!}$ is isomorphic to the image of $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ under $e_{\pi(\phi)}$. Thus, we will be able to consider $H^{2}(\phi)_{!}$ both as a subobject and as a quotient of the Hodge structure on $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$.
@@
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the assumption of Prop. 6.8, but $\phi$ is of infinity type $(-k-3,k)$. The same proof, but employing the local system $V^{\vee}_{k}$ instead of $V_{k}$, shows the existence of a Hodge structure $H^{2}(\phi)_{!}$ of type $(0,k+2)$ satisfying all the other requirements in the statement. This follows from the computation of the highest weight of $V^{\vee}_{k}$ of Rmk. 2.2 and from the fact that using the latter local system, by Rmk. 6.7, the infinite component $\pi(\phi)_{\infty}$ is isomorphic to the discrete representation $\pi^{-}$, which is anti-holomorphic by ([7, 1.3]).
7. Extensions and biextensions associated to Hecke characters
In this last section, we keep on considering Hecke characters $\phi$ of $F$ satisfying the hypotheses of the previous section, which will be recalled below. In particular, their $L$-function is supposed to vanish at the central point, and as a consequence, as discussed in 3.3, the Bloch-Beilinson conjectures predict the existence of a non-trivial extension of Hodge structures $E_{\phi}$ of geometric origin.
Our first aim is to put together all the ingredients introduced so far, in order to construct candidates $E_{\phi}(\Psi)$ for $E_{\phi}$, by means of the cohomology of a local system on a Picard modular surface $\mathrm{S}_{K}$. They depend on the choice of a cohomology class $\Psi$ in a suitable subspace of Eisenstein cohomology of $\mathrm{S}_{K}$. This will prove Theorem 1 in the introduction.
In the second and last part of this section, we will show how such $E_{\phi}(\Psi)$’s sit naturally in certain biextensions of Hodge structures. The relevance of this construction for our program of proving non-triviality of (some) $E_{\phi}(\Psi)$ has been discussed in the introduction.
@@ 7.1. Construction of extensions
In order to state our main theorem, recall that to a Hecke character $\phi$ satisfying the hypotheses formulated below, we have attached pure Hodge structures $\partial H^{2}(\phi)$ and $H(\phi)_{!}$, respectively in (43) and (54).
@@
Let $k$ be a positive integer and $\phi$ an Hecke character of $F$ of type $(k,-k-3)$, satisfying
@@
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type $(k,-(k+3))$ attached to $\phi$.
@@
    Since $\epsilon(\phi)=-1$, we have $L(\phi,-1)=0$. Thus, by Corollary 5.15, the pure Hodge structure $\partial H^{2}(\phi)$ is a sub-Hodge structure of $H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. By Proposition 6.8 and Remark 6.9, the pure Hodge structure $H^{2}(\phi)_{!}$ is a quotient Hodge structure of $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. Hence, the desired extension is obtained from the short exact sequence (55) by taking pullback via the inclusion $\partial H^{2}(\phi)\hookrightarrow H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ and pushout via the projection $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})\twoheadrightarrow H^{2}(\phi)_{!}$.
@@
    Since the type of $H^{2}(\phi)_{!}$ is $(k+2,0)$ and the type of $I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is $(1,k+2)$, the type of the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is $(k+1,-k-2)$. This tells us that as a $\overline{{\mathbb{Q}}}$-Hodge structure, $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is indeed isomorphic to $H_{\phi}(-1)$.
@@
Let $k$ and $\phi$ be as in Theorem 57. The choice of an element $\Psi$ of $I_{\phi}^{K}$ provides, by pullback of the extension (57), a 1-extension $E_{\phi}(\Psi)$ of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
@@
Using Remarks 5.18 and 6.10, one can run the same proof as the one of Theorem 57, but employing $V^{\vee}_{k}$ instead of $V_{k}$, and deduce the following variant of the above Corollary.
@@
Let $k$ and $\phi_{{\mathbb{Q}}}$ be as in Theorem 57, but suppose that $\phi$ is of infinity type $(-k-3,k)$. The choice of an element $\Psi$ of $I_{\phi}^{K}$ provides a 1-extension $E_{\phi}(\Psi)$ of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
@@
For the needs of the next subsection, we also need to construct a dual extension, in the following way. The extension of scalars to $\overline{{\mathbb{Q}}}$ of the long exact sequence (28) gives rise to a short exact sequence
@@
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type $(k,-(k+3))$ attached to $\phi$.
@@
[…]
@@
7.2. Construction of biextensions
Consider the extensions $E_{\phi}(\Psi)$ constructed in the previous subsection. To verify the Bloch-Beilinson conjectures for our Hecke characters $\phi$, one has to show that there exists an element $\Psi$ of $I^{K}_{\phi}$ such that the extension $E_{\phi}(\Psi)$ is non trivial. This is the subject of work in progress by the authors, where verifying the desired non-triviality will rest crucially on the non-vanishing of a certain biextension height ([15], [9]) associated to $E_{\phi}(\Psi)$. Let us then define the input for such a height, i.e. a biextension in which $E_{\phi}(\Psi)$ sits.
We keep on looking at the local system $V_{k}$ on the Picard modular surface $\mathrm{S}_{K}$. As in subsection 4.3, we consider the Baily-Borel compactification $\mathrm{S}^{*}_{K}$ together with the open immersion $j$ of $\mathrm{S}_{K}$ and the closed immersion $i$ of the boundary $\partial\mathrm{S}_{K}$. To simplify notation, we will denote the latter by $\partial$.
@@ 7.2. Construction of biextensions
For any $\triangle\in\{\Theta,\Sigma\}$, consider the pair of complementary, open resp. closed immersions
@@
${j_{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${i_{*}i^{*}j_{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\triangle}}$${i_{\triangle,*}i^{\triangle,*}j^{\triangle}_{*}}$
@@
by restriction of those defined in (43), resp. (44).
@@
(67) $0\rightarrow H^{2}(\phi)_{!}\rightarrow\tilde{E}\rightarrow\partial H_{\Theta}^{2}(\phi)\rightarrow 0$
@@
$H^{1}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\twoheadrightarrow I^{K}_{\theta\phi,\Sigma}$
@@
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{\tilde{E}}_{\Sigma}\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
@@
(70) $0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}_{\Sigma}\rightarrow\tilde{\tilde{E}}\rightarrow 0$
An analogous pullback applied to the extension (56) yields an extension $\tilde{E}$ as in in (67). Pullback of (70) along the induced morphism
@@
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}^{\sharp}\rightarrow\tilde{E}\rightarrow 0$
which gives the right four-term column of the diagram in the statement. ∎
@@
$\partial=\Theta\bigsqcup\Sigma\,.$
@@
To conclude our construction, we need to recall the concept of biextension (cfr. for example [9], Def. 2.1) in our setting.
Definition 7.10.
Let $H$ be a pure $\overline{{\mathbb{Q}}}$-Hodge structure of weight $-1$. A biextension of $H$ is a $\overline{{\mathbb{Q}}}$-mixed Hodge structure $B$ together with an identification of its associated weight-graded semisimple Hodge structure
$\mathop{\rm Gr}\nolimits^{W}B\simeq\mathds{1}(1)\oplus H\oplus\mathds{1}$
@@
$\partial=\Theta\bigsqcup\Sigma\,.$
@@
the diagram of Corollary 7.9 induces a biextension

### v4

@@ 6.2. An application of Rogawski’s results
    is the same as the dimension of the space $V^{2}(\pi_{f}))$ of Case 5 of [37, pp. 91-92], which according to loc. cit. is equal to 1, by our choice of the archimedean component of $\pi(\phi)$. We then obtain the desired Hecke submodule. Finally, to get the assertion on the type of $H^{2}(\phi)_{!}$, we apply the following two facts:
    (a) if $r$ is the weight of the pure Hodge structure on $V_{\lambda}$ (so that, by (23), $r=k$ in our case), then $H^{2}_{!}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ has weight $r+2$. Thus, its Hodge substructure $H^{2}(\pi(\phi)_{f})$ has the same weight;
    (b) the Hodge types of $H^{2}(\pi(\phi)_{f})$ are those $(p,q)$’s appearing in the $(p,q)$-decomposition of the space $H^{2}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi(\phi)_{\infty}\otimes V_{k,{\mathbb{C}}})$, which is in turn determined by the $(p,q)$-decomposition of the discrete series representation $\pi(\phi)_{\infty}$. Now, by Remark 6.7 the discrete series representation $\pi(\phi)_{\infty}$ coincides with $\pi^{+}$, which is holomorphic by ([10, Sect. 1.3]).
@@
Suppose the hypotheses of Proposition 6.8 hold. For any small enough level $K$, the simple, level $K$ Hecke module $\pi^{K}(\phi)_{f}$ corresponds to a minimal ideal in the level $K$ Hecke algebra. Take an idempotent generator $e_{\pi(\phi)}$ of this ideal. By a standard argument (cfr. for example [42, Prop. 5.4]) we see that the Hodge structure $H^{2}(\phi)_{!}$ is isomorphic to the image of $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ under $e_{\pi(\phi)}$. Thus, we will be able to consider $H^{2}(\phi)_{!}$ both as a subobject and as a quotient of the Hodge structure on $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$.
@@
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the assumption of Proposition 6.8, but with $\phi$ being of infinity type (k3,k)(-k-3,k) instead of (k,k3)(k,-k-3). The same proof, but employing the local system $V^{\vee}_{k}$ instead of $V_{k}$, shows the existence of a Hodge structure $H^{2}(\phi)_{!}$ of type $(0,k+2)$ satisfying all the other requirements in the statement. This follows from the computation of the highest weight of $V^{\vee}_{k}$ of Remark 2.2 and from the fact that using the latter local system, by Remark 6.7, the component at infinity $\pi(\phi)_{\infty}$ is isomorphic to the discrete representation $\pi^{-}$, which is anti-holomorphic by ([10, 1.3]).
7. Extensions associated to Hecke characters
In this last section, we continue considering Hecke characters $\phi$ of $F$ satisfying the hypotheses of the previous section, which will be recalled below. In particular, their $L$-function is supposed to vanish at the central point, and as a consequence, as discussed in 3.4, the Bloch-Beilinson conjectures predict the existence of a non-trivial extension of Hodge structures $E_{\phi}$ of geometric origin.
Our first aim is to put together all the ingredients introduced so far, to construct candidates $E_{\phi}(\Psi)$ for $E_{\phi}$, by means of the cohomology of a local system on a Picard modular surface $\mathrm{S}_{K}$. They depend on the choice of a cohomology class $\Psi$ in a suitable subspace of Eisenstein cohomology of $\mathrm{S}_{K}$. This will prove Theorem 1 stated in the introduction.
@@ 7.1. Construction of extensions
To state our main theorem, recall that to a Hecke character $\phi$ satisfying the hypotheses formulated below, we have attached pure Hodge structures $\partial H^{2}(\phi)$ and $H^{2}(\phi)_{!}$, mentioned respectively in (48) and (61).
@@
Let $k$ be a positive integer and $\phi$ a Hecke character of $F$ of type (k,k3)(k,-k-3), satisfying
@@
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type (k,k3)(k,-k-3) attached to $\phi$.
  3. (3)
    Any element $\Psi$ of $I_{\phi}^{K}$ provides an extension $E_{\phi}(\Psi)$ of geometric origin of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$.
@@
    We know that the hypothesis on $\phi_{{\mathbb{Q}}}$ is equivalent to asking that $\phi$ verifies (53). By Remark 3.2, there is a well-defined sign $\epsilon(\phi)$, and $\epsilon(\phi)=-1$ implies $L(\phi,-1)=0$. Thus, by Corollary 5.15, the pure Hodge structure $\partial H^{2}(\phi)$ is a sub-Hodge structure of $H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. By Proposition 6.8 and Remark 6.9, the pure Hodge structure $H^{2}(\phi)_{!}$ is a quotient Hodge structure of $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. Hence, the desired extension is obtained from the short exact sequence (62) by taking pullback via the inclusion $\partial H^{2}(\phi)\hookrightarrow H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ and pushout via the projection $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})\twoheadrightarrow H^{2}(\phi)_{!}$.
@@
    Since the type of $H^{2}(\phi)_{!}$ is $(k+2,0)$ and the type of $I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is $(1,k+2)$, the type of the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is (k+1,k2)(k+1,-k-2). This tells us that as a $\overline{{\mathbb{Q}}}$-Hodge structure, $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is indeed isomorphic to $H_{\phi}(-1)$.
  3. (3)
    Choosing an element $\Psi$ of $I_{\phi}^{K}$ amounts to defining a morphism of Hodge structures $\mathds{1}\rightarrow\mathds{1}^{m(\phi,K)}$. By pullback of the extension (64) along this morphism, we get the desired extension $E_{\phi}(\Psi)$. To see that it is of geometric origin, denote by $M(V_{k})$ the motive $e_{\lambda}M(A^{N})$ obtained as in (31) when choosing the representation $V_{\lambda}$ to be equal to $V_{k}$. It has Hodge realization $R_{H}(M(V_{k}))=H^{\bullet}(\mathrm{S}_{K},V_{k})$. Then, by construction, $E_{\phi}(\Psi)$ is a subquotient Hodge structure of the mixed Hodge structure $R_{H}(M(V_{k}))$, hence it is of geometric origin.
@@
Let $k$ and $\phi$ be as in Theorem 7.1. Associating the extension $E_{\phi}(\Psi)$ to $\Psi\in I_{\phi}^{K}$ defines a morphism
@@
Using Remarks 5.18 and 6.10, one can run the same proof as the one of Theorem 7.1, but employing $V^{\vee}_{k}$ instead of $V_{k}$, and deduce the following variant of point (3) of the above Theorem.
@@
Let $k$ and $\phi_{{\mathbb{Q}}}$ be as in Theorem 7.1, but suppose that $\phi$ is of infinity type (k3,k)(-k-3,k). The choice of an element $\Psi$ of $I_{\phi}^{K}$ provides an extension $E_{\phi}(\Psi)$ of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
@@
such that all extensions in its image are of geometric origin.
For the needs of the next subsection, we also need to construct a dual extension in the following way. The extension of scalars to $\overline{{\mathbb{Q}}}$ of the long exact sequence (33) gives rise to a short exact sequence
@@
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type (k,(k+3))(k,-(k+3)) attached to $\phi$.
@@
The choice of an element $\Psi^{\prime}$ of $I_{\theta\phi}^{K}$ provides, by pushout of the extension (69), an extension $E_{\phi}(\Psi^{\prime})$ of $H_{\phi}(-1)$ by $\mathds{1}(1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
@@
such that all extensions in its image are of geometric origin.
8. Biextensions associated to Hecke characters
Consider the extensions of geometric origin $E_{\phi}(\Psi)$ constructed in the previous section. To verify the Bloch-Beilinson conjectures for our Hecke characters $\phi$, one has to show that there exists an element $\Psi$ of $I^{K}_{\phi}$ such that the extension $E_{\phi}(\Psi)$ is non-trivial. This is the subject of work in progress by the authors, where, as discussed in the introduction, verifying the desired non-triviality will rest crucially on the non-vanishing of a certain biextension height ([21], [12]) associated to $E_{\phi}(\Psi)$.
The aim of this section is then to define the input for such a height, that is, a biextension in which $E_{\phi}(\Psi)$ sits. First, we recall what this means in our setting (cfr. for example [12], Def. 2.1).
Definition 8.1.
Let $H$ be a pure $\overline{{\mathbb{Q}}}$-Hodge structure of weight 1-1. A biextension of $H$ is a $\overline{{\mathbb{Q}}}$-mixed Hodge structure $B$ together with an identification of its associated weight-graded semisimple Hodge structure
$\mathop{\rm Gr}\nolimits^{W}B\simeq\mathds{1}(1)\oplus H\oplus\mathds{1}\,.$
Now, we continue to look at the local system $V_{k}$ on the Picard modular surface $\mathrm{S}_{K}$. As in Subsection 4.3, we consider the Baily-Borel compactification $\mathrm{S}^{*}_{K}$ together with the open immersion $j$ of $\mathrm{S}_{K}$ and the closed immersion $i$ of the boundary $\partial\mathrm{S}_{K}$. To simplify notation, we will denote the latter by $\partial$. In the following, we make free use of the six functor formalism for mixed Hodge modules mentioned in Subsection 2.9. We will work in the derived category of mixed Hodge modules, with its canonical triangulated structure; all functors will be implicitly derived (e.g., we denote $Rj_{*}$ simply by $j_{*}$), and we will not use different symbols for cohomology and hypercohomology.
@@ 8. Biextensions associated to Hecke characters
For any $\triangle\in\{\Theta,\Sigma\}$, consider the pair of complementary, open respectively closed immersions
@@
[…]
@@
by restriction of those defined in (48), respectively (49).
@@
(74) $0\rightarrow H^{2}(\phi)_{!}\rightarrow\tilde{E}\rightarrow\partial H_{\Theta}^{2}(\phi)\rightarrow 0\,$,
@@
$H^{1}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\twoheadrightarrow I^{K}_{\theta\phi,\Sigma}\,.$
@@
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{\tilde{E}}_{\Sigma}\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\,.$
@@
(77) $0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}_{\Sigma}\rightarrow\tilde{\tilde{E}}\rightarrow 0\,.$
An analogous pullback applied to the extension (63) yields an extension $\tilde{E}$ as in (74). Pullback of (77) along the induced morphism
@@
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}^{\sharp}\rightarrow\tilde{E}\rightarrow 0\,.$
By construction, this extension fits in the diagram given in the statement, as its right four-term column. ∎
@@
$\partial=\Theta\bigsqcup\Sigma$
@@
$\partial=\Theta\bigsqcup\Sigma\,$
@@
the diagram of Corollary 8.4 induces a biextension of geometric origin